Hi!
I needed help understanding the reasoning behind each step in this solution/problem...
33 1/3% = 0.33 1/3 = 33 1/3 /100 = 100/3 / 100/1 = 100/3 * 1/100 = 1/3
Particularly why 1/3 is not changed in the first few steps.

Answer by drk(1908)

33 1/3% = 0.33 1/3
we are changing a percent to a decimal. Percent means number / 100, so just move the decimal 2 places left to make this conversion.
0.33 1/3 = (33 1/3) / 100
here we are changing the decimal to a fraction. .33 means 33/100.
(33 1/3) /100 = (100/3) / (100/1)
here w are urning mixed and proper fractions to improper fractions. 33 1/3 = 100/3
(100/3) / (100/1) = 100/3 * 1/100
here we are converting from division to multiplication. In order to do this we flip the second fraction.
100/3 * 1/100 = 1/3
here we are multiplying the numerators and then denominators together to get 100/300 which reduces to 1/3